can you hlep me

Question 6 Let r = xi+yj + zk and r = r. Find V.r. O A.3 B. O C.1 D. none of these O E.4
Answer #1

Giron Ju= ait yj+Zk Consider you = lijf+*).(ai +43 +3K) -K (2) + = (4)+ (3) = 1+1+1 3. Hence to N = 3 : 3 Correct choice is a